You have plenty of cash to invest. You are considering an investment of $125,000 in a...

  1. You have plenty of cash to invest. You are considering an investment of $125,000 in a project which is expected to earn $14425 a year for ten years. Is it an attractive investment if your minimum expected annual rate of return is 5% (compound interest)? Calculate the expected annual rate of return (R) of the project. (6 points) (Please give proper steps by using the Present Value of an Annuity!
Answer #1

Let IRR be i%, as PW at IRR = 0

-125000 + 14425 * (P/A,i%,10) = 0

(P/A,i%,10) = 125000 / 14425 = 8.665511

using trail and error method

When i = 5%, value of (P/A,i%,10) = 7.721735

When i = 3%, value of (P/A,i%,10) = 8.530203

When i = 2%, value of (P/A,i%,10) = 8.982585

using interpolation

i = 2% + (8.982585 - 8.665511) / (8.982585 - 8.530203)*(3% - 2%)

i = 2% + 0.700898%

i = 2.7% (Approx)

As IRR < MARR, this is not a good invesment

Pls comment if you think final solution is wrong or you require further explanation